Circular No. 8668

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 SB_{216} (LONEOS) An apparently asteroidal object discovered by the LONEOS project

that was designated 2005 SB_{216} by the Minor Planet Center (discovery position below from MPS 144403) has been found to show cometary appearance. S. Foglia, Novara Veveri, Italy, reports that CCD observations that he obtained with G. Galli, S. Minuto, and D. Crespi using a 0.40-m f/4 reflector on 2005 Dec. 7 and 29 show the object to be slightly diffuse; Foglia adds that CCD frames taken by L. Buzzi with a 0.60-m f/4.6 reflector at Varese on 2006 Feb. 4.8 UT also show diffuseness. F. Bernardi, D. J. Tholen, and J. Pittichová, University of Hawaii (UH), report that two R-band 600-s exposures taken with the UH 2.2-m reflector at Mauna Kea on Feb. 7.27 show the object to be fuzzy and slightly elongated ∼ 2′′.7 in p.a. ∼ 78^{o}; in an aperture of 4′′.4, the magnitude was 19.2.

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.30.38046  0^{h}54^{m}39.^{s}28  +14^{o}22′54′′.3  19.1

The following orbital elements are taken from MPEC 2006-C48:

Epoch = 2007 Mar. 1.0 TT

T = 2007 Feb. 11.3910 TT  ω = 83.^{o}5889
e = 0.463528  Ω = 1.6983 }2000.0
q = 3.817862 AU  i = 24.0974

a = 7.116613 AU n^{o} = 0.0519151 P = 18.98 years

SUPERNOVAE 2006Y AND 2006Z Two apparent supernovae have been reported: 2006Y by P. Luckas, O.

Trondal, and M. Schwartz (cf. IAUC 8655; unfiltered CCD frames, 0.35-m Tenagra telescope at Perth) and 2006Z by the SDSS collaboration (found in spectra; communicated by M. SubbaRao, University of Chicago and Adler Planetarium; cf. IAUC 8513; r-band magnitude given below).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006Y Feb. 3.58 7^{h}13^{m}17.^{s}19 −51^{o}41′18′′.8 17.7 1′′.7 W, 5′′.2 N
2006Z Feb. 1.51 13 44 58.07 +26 18 25.7 20.4  —

Additional unfiltered magnitudes of 2006Y: Jan. 27.59 UT, [18.5; Feb. 7.60, 17.3. SN 2006Z, which appears to be a type-Ia supernova with an age of ≈ −7 ± 4 days after maximum light, is coincident with the center of the host galaxy (r = 17.0 from an image taken on 2004 June 12; z = 0.1232).

2006 February 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green